can you help me find an equation in slop-intercept from that passes through the given point and is parallel to the given equation.?
Well:
1.(5,2), y=(4/5)x+2
2.(-3,5), 3y+6x=-18
3.(2,-2), 2y=3x+6

Answer by mananth(16946)

1 y = 4/5 x + 2
Divide by 1
y = 4/5 x + 2
Compare this equation with y=mx+b
slope m = 4/5
The slope of a line parallel to the above line will be the same
The slope of the required line will be 4/5
m= 4/5 ,point (5,2)
Find b by plugging the values of m & the point in
y=mx+b
2=4+ b
b= 2
m= 4/5
Plug value of the slope and b in y = mx +b
The required equation is y = 4/5 x + 2
-----
6 x + 3 y = -18
Find the slope of this line
make y the subject
3 y = -6 x + -18
Divide by 3
y = -2x-6
Compare this equation with y=mx+b
slope m = -2
The slope of a line parallel to the above line will be the same
The slope of the required line will be -2
m= -2 ,point ( -3 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = 6 + b
b= 1
m= -2
Plug value of the slope and b in y = mx +b
The required equation is y=-2x+1
similarly you can work out the third one

m.ananth@hotmail.ca




m.ananth@hotmail.ca